OpenStudy (anonymous):

How would you graph this - 5 = 2/3 (x - 4).

OpenStudy (anonymous):

Solve for x: -5=2/3(x-4) -15/2=x-4 -7/2 =x find where x=-7/2 on a graph

OpenStudy (anonymous):

How would that look on a graph?